Show on the dimensional Analysis that v^2 - u^2 = 2as where u is the initial velocity, v is the final velocity, a is the acceleration due to gravity and s is the distance moved

1
Expert's answer
2022-01-10T09:10:28-0500

Perform the dimensional analysis:


"[v]=[m/s],\\
[a]=[m/s^2],\\
[d]=[m].\\\space\\
[m/s]^2=[m/s^2]·m,\\
m^2/s^2=m·m/s^2,\\
m^2/s^2=m^2/s^2."